# Xinxu Weia    Kanhao Zhaoa    Yong Jiaoa    Lifang Hea    Yu Zhanga11footnotemark: 1

### v2

@@ Abstract
As large language models (LLMs) continue to revolutionize AI research, there is a growing interest in building large-scale brain foundation models to advance neuroscience. While most existing brain foundation models are pre-trained on time-series signals or connectome features, we propose a novel graph-based pre-training paradigm for constructing a brain graph foundation model. In this paper, we introduce the Brain Graph Foundation Model, termed BrainGFM, a unified framework that leverages graph contrastive learning and graph masked autoencoders for large-scale fMRI-based pre-training. BrainGFM is pre-trained on a diverse mixture of brain atlases with varying parcellations, significantly expanding the pre-training corpus and enhancing the model’s ability to generalize across heterogeneous fMRI-derived brain representations. To support efficient and versatile downstream transfer, we integrate both graph prompts and language prompts into the model design, enabling BrainGFM to flexibly adapt to a wide range of atlases, neurological and psychiatric disorders, and task settings. Furthermore, we employ meta-learning to optimize the graph prompts, facilitating strong generalization to previously unseen disorders under both few-shot and zero-shot learning conditions via language-guided prompting. BrainGFM is pre-trained on 27 neuroimaging datasets spanning 25 common neurological and psychiatric disorders, encompassing 2 types of brain atlases (functional and anatomical) across 8 widely-used parcellations, and covering over 25,000 subjects, 60,000 fMRI scans, and a total of 400,000 graph samples aggregated across all atlases and parcellations.
@@ 1 Introduction
With the rise of large language models (LLMs) (Achiam et al., 2023), large-scale pre-trained foundation models (FMs) have been proposed across various domains, including computer vision (Touvron et al., 2023), natural language processing (Achiam et al., 2023), and data mining (Xia et al., 2024). Recently, the field of neuroscience has also begun to witness the emergence of brain foundation models. As a widely used data modality in neuroscience, functional magnetic resonance imaging (fMRI) (Markiewicz et al., 2021; Bycroft et al., 2018) plays a crucial role in understanding brain function and dysfunction. Developing a fMRI-based brain foundation model is of great importance for advancing neuroscience and its translational research. Due to the high complexity and cost of fMRI data acquisition (Van Essen et al., 2012; Cui et al., 2022), coupled with strong heterogeneity and substantial inter-subject variability, most existing traditional deep learning-based fMRI models (Wei et al., 2025; Kan et al., 2022; Jiao et al., 2025) are trained on relatively small datasets. Consequently, these models are typically tailored to specific tasks, disorders, or cohorts, resulting in limited generalizability, poor flexibility, and weak transferability to unseen tasks, datasets or disorder conditions. Furthermore, training on small-scale datasets often leads to under-fitting, ultimately compromising model performance and reliability. These issues have become common limitations of traditional deep learning models for fMRI. However, they can be effectively addressed by building fMRI brain FMs. Since FMs are typically pre-trained on large-scale datasets with rich diversity (Touvron et al., 2023; Xia et al., 2024), spanning various data types and knowledge representations within the neuroscience domain, the resulting models exhibit strong generalization and broad applicability.
@@ 2.3 Meta-Learning
Figure 1: The pipeline of our proposed BrainGFM. (a) A large-scale brain fMRI graph dataset is constructed for pre-training. (b) BrainGFM is pre-trained using graph contrastive and masked autoencoder strategies, with atlas/parcellation tokens [A/P] to encode atlas-specific information. (c) We introduce graph prompts and use meta-learning to optimize them for few-shot adaptation, keeping the graph FM backbone frozen. (d) Finally, we freeze both the model and graph prompts, and use language prompts to enable zero-shot transfer to new tasks. Note that "Schf." means Schaefer atlas.
@@ 3.2 Graph Pre-Training for Building BrainGFM
We adopt a Graph Transformer (Yun et al., 2019) as the backbone of our BrainGFM. As illustrated in Figure 1(b), we first transform the input fMRI brain graphs and project them to obtain brain graph embeddings, where each token corresponds to a brain ROI. We apply Positional Encoding Tokens, denoted as [PE], to each brain ROI, enabling the model to perceive and learn the topological and spatial characteristics of each ROI. Unlike conventional positional encodings used in standard Transformer models (Vaswani et al., 2017), graph-based positional encodings are inherently different, as they require encoding the relative positions between nodes in the graph structure. Compared to the commonly used Laplacian positional encoding (Dwivedi et al., 2023) and node degree positional encoding (You et al., 2019) in graph-based models, we adopt a more efficient alternative Random Walk Structural Encoding (RWSE) (Dwivedi et al., 2021) as our positional encoding strategy. More details about PEs can be found in Appendix I. Furthermore, inspired by language models in NLP, we insert Atlas/Parcellation Tokens, denoted as [A/P], to the brain graph embeddings during pre-training, enabling the model to better distinguish and learn from different atlases and parcellations. Note that the construction of [A/P] tokens is described in detail in Section 3.4. Incorporating this token enables the model to capture parcellation-specific patterns, which is crucial as prior studies (Hermosillo et al., 2024; Liu et al., 2023; Wu et al., 2025) show that different brain disorders are better represented by specific parcellations. For instance, MDD benefits from Schaefer200 or Power264 and ASD is better captured by Shen268 or Schaefer200. Embedding such parcellation-aware information helps improve model generalization across disorders and atlas configurations.
@@ 3.4 Zero-Shot Graph/Language Prompt-Tuning
Building on the few-shot capability, we further introduce language prompts to enable more generalized zero-shot learning by jointly guiding both the graph prompt and the pre-trained foundation model. In zero-shot scenarios, the parameters of the graph prompt are also frozen, meaning the model cannot rely on prompt adaptation through learning. Instead, the language prompt provides semantic guidance, allowing the model to generalize and adapt to unseen downstream data, tasks, and disorder types without any gradient-based updates. As shown in Figure 1(d), in order to enable the model to recognize and distinguish between different tasks and disorders in zero-shot settings, we introduce Task/Disorder Tokens, denoted as [T/D], during the downstream fine-tuning stage, following a similar design with Atlas/Parcellation Tokens [A/P]. To construct the [T/D] tokens, we first generate a textual description for each disorder, including its full name, abbreviation, and a concise clinical summary. For example, for Major Depressive Disorder, the corresponding text description is: Major Depressive Disorder (MDD) is a common mental illness characterized by persistent and profound low mood, loss of interest, and cognitive impairment, significantly affecting daily life and social functioning.” (Otte et al., 2016) We then encode these textual descriptions using a BERT model (Devlin et al., 2019) pre-trained on large-scale medical corpora, such as BioClinicalBERT (Huang et al., 2019; Alsentzer et al., 2019), to obtain semantic-rich text embeddings. These embeddings are subsequently projected and embedded as [T/D] tokens, which are incorporated into the model during downstream adaptation. Similarly, the construction of [A/P] tokens is also based on language text. For each atlas and parcellation, we provide a textual description of its name, such as “Schaefer100”, “Schaefer200”, or “AAL116”. These text descriptions are then encoded using the BioClinicalBERT pre-trained model to extract language embeddings, which are subsequently transformed into [A/P] tokens. As shown in Figure 1(d), the [T/D] and [A/P] tokens are ultimately concatenated with the ROI tokens from the graph embeddings as language prompt tokens. This combined input is then fed into the foundation model to guide feature extraction specific to the given dataset, task, and disorder. By introducing disorder-specific semantic priors through the [T/D] and [A/P] tokens, the model is better equipped to capture characteristics from different tasks, disorders, atlas and parcellations, thereby improving its downstream adaptation ability in zero-shot settings without any training.
@@ 4.1 Comparison with Other Methods
BrainLM 67.6±1.9 69.1±2.2 64.0±2.7 71.7±2.0 68.1±2.6 71.6±1.8 66.9±2.5 69.6±2.4 78.3±1.7 82.9±2.5 73.1±2.6 81.4±2.7 79.3±2.0 83.2±2.1 83.5±2.8 75.8±2.4 82.6±2.5 85.0±2.6 78.6±2.0 86.5±2.7
BrainMass 67.0±2.3 67.5±2.0 64.7±2.8 71.1±1.8 68.9±2.1 70.1±1.9 69.5±2.4 66.3±2.6 77.8±2.3 82.7±2.7 72.6±2.9 81.4±2.5 76.9±2.4 80.2±2.8 83.6±2.5 76.1±2.6 81.0±2.7 84.0±2.2 80.3±2.3 81.9±2.9
Brain-JEPA 69.8±1.9 71.6±2.0 66.2±1.6 72.9±2.4 70.1±1.7 73.8±1.8 70.1±1.9 69.3±1.7 79.1±2.2 84.3±1.8 76.8±2.2 83.6±2.1 83.4±1.7 85.9±1.3 84.3±1.9 77.5±1.5 85.4±1.8 86.7±2.0 85.3±2.2 83.7±1.9
BrainGFM 70.6±1.6 72.2±1.5 67.3±1.7 73.4±2.1 71.2±1.9 73.5±1.4 70.4±1.5 69.8±1.7 80.3±2.6 85.1±2.2 76.2±1.5 84.4±1.7 83.6±1.6 85.5±1.7 85.8±1.6 77.9±1.9 85.2±1.6 86.3±2.1 87.7±1.9 82.6±1.7
@@ 4.2 Ablation Study for Full/Few/Zero-Shot on Graph/Language Prompt and Meta Learning
4.3 Experiments on Different Atlases and Parcellations with Different ROI Resolutions
@@ 4.4 Comparison Among Time-Series, ROI and Graph-Based Foundation Models
Figure 4: Comparison of performance and efficiency across different brain FMs.
@@ 4.6 Ablation Study on Different Foundation Model Pre-Training Methods
As shown in Figure 5, we compare the effectiveness of different graph pre-training strategies, including graph contrastive learning (GCL), graph-masked autoencoders (GMAE) and their sequential combination. The results demonstrate that GCL slightly outperforms GMAE, and that combining GCL and GMAE yields further performance gains compared to using either method alone.
@@
Graph contrastive learning (GCL) pre-training primarily focuses on capturing global representations of brain graphs by encouraging the model to aggregate holistic graph-level features and distinguish between different graph attributes and categories. In comparison, graph masked autoencoders (GMAE) pre-training emphasizes the learning of local representations, where the model reconstructs masked brain ROIs based on information from their local neighborhoods, thereby promoting specialization in ROI-level feature extraction. By sequentially combining GCL and GMAE during pre-training, BrainGFM is able to simultaneously acquire both global and local discriminative capabilities. Notably, the integration of global and local information has been widely recognized as critical for understanding brain organization and pathology in neuroscience and neuroimaging studies. Consequently, our pre-trained model benefits from this multi-scale representation learning, leading to enhanced transferability and improved performance across various downstream tasks.
@@ Appendix E Ablation Study for Different Brain Graph Construction Methods
[…]
Corpus Top-K Correlation KNN ABIDE II ASD 73.5 73.9 HBN MDD 85.5 86.7
@@ Appendix G Ablation Study on Different Tuning Methods
As shown in Table 5, after completing model pre-training, we explored three downstream adaptation strategies: full fine-tuning [Sun et al., 2022], parameter-efficient fine-tuning (PEFT) [Ding et al., 2023], and graph prompt-tuning [Sun et al., 2023]. Full fine-tuning updates all model parameters during downstream training, offering strong performance but at a high computational cost. In contrast, PEFT methods reduce training overhead by modifying only a small subset of parameters or introducing lightweight modules. We specifically evaluate two popular PEFT variants: prefix-tuning and LoRA. Note that the details of all tuning methods are summarized in Table 6.
Graph prompt-tuning further improves efficiency by freezing the entire pre-trained model and updating only a small set of learnable prompt vectors injected into the input. This strategy allows the model to adapt without altering its core parameters, making it highly suitable for resource-constrained settings. In full-shot scenarios, fine-tuning delivers the best results, but PEFT methods achieve competitive performance with significantly lower computational demands. Graph prompt-tuning, while slightly less accurate, offers the best efficiency–adaptability trade-off by minimizing trainable parameters. Given the structural nature of brain graphs, where node and edge features capture complex spatial and relational dependencies, we examine two prompt insertion mechanisms: addition (“+”) and multiplication (“$\times$”). Results show that multiplicative insertion consistently outperforms the additive version, likely because scaling features better preserves relational patterns. Moreover, to further leverage topological information, we extend the prompt-tuning framework by incorporating edge prompts in addition to node prompts. This design grants the model greater flexibility to adjust local connectivity, leading to improved transfer performance across diverse downstream tasks.

### v3

# A Brain Graph Foundation Model:Pre-Training and Prompt-Tuning across Broad Atlases and Disorders

@@ Abstract
As large language models (LLMs) continue to revolutionize AI research, there is a growing interest in building large-scale brain foundation models to advance neuroscience. While most existing brain foundation models are pre-trained on time-series signals or connectome features, we propose a novel graph-based pre-training paradigm for constructing a brain graph foundation model. In this paper, we introduce the Brain Graph Foundation Model, termed BrainGFM, a unified framework that leverages graph contrastive learning and graph masked autoencoders for large-scale fMRI-based pre-training. BrainGFM is pre-trained on a diverse mixture of brain atlases with varying parcellations, significantly expanding the pre-training corpus and enhancing the model’s ability to generalize across heterogeneous fMRI-derived brain representations. To support efficient and versatile downstream transfer, we integrate both graph prompts and language prompts into the model design, enabling BrainGFM to flexibly adapt to a wide range of atlases, neurological and psychiatric disorders, and task settings. Furthermore, we employ meta-learning to optimize the graph prompts, facilitating strong generalization to previously unseen disorders under both few-shot and zero-shot learning conditions via language-guided prompting. BrainGFM is established on 27 neuroimaging datasets spanning 25 common neurological and psychiatric disorders, encompassing 2 types of brain atlases (functional and anatomical) across 8 widely-used parcellations, and covering over 25,000 subjects, 60,000 fMRI scans, and a total of 400,000 graph samples aggregated across all atlases and parcellations.
@@ 1 Introduction
With the rise of large language models (Achiam et al., 2023), large-scale pre-trained foundation models (FMs) have been proposed across various domains, including computer vision (Touvron et al., 2023), natural language processing (Achiam et al., 2023), and data mining (Xia et al., 2024). Recently, the field of neuroscience has also begun to witness the emergence of brain foundation models. As a widely used data modality in neuroscience, functional magnetic resonance imaging (fMRI) (Markiewicz et al., 2021; Bycroft et al., 2018) plays a crucial role in understanding brain function and dysfunction. Developing a fMRI-based brain foundation model is of great importance for advancing neuroscience and its translational research. Due to the high complexity and cost of fMRI data acquisition (Van Essen et al., 2012; Cui et al., 2022), coupled with strong heterogeneity and substantial inter-subject variability, most existing traditional deep learning-based fMRI models (Wei et al., 2025; Kan et al., 2022; Jiao et al., 2025) are trained on relatively small datasets. Consequently, these models are typically tailored to specific tasks, disorders, or cohorts, resulting in limited generalizability, poor flexibility, and weak transferability to unseen tasks, datasets or disorder conditions. Furthermore, training on small-scale datasets often leads to under-fitting, ultimately compromising model performance and reliability. These issues have become common limitations of traditional deep learning models for fMRI. However, they can be effectively addressed by building fMRI brain FMs. Since FMs are typically pre-trained on large-scale datasets with rich diversity (Touvron et al., 2023; Xia et al., 2024), spanning various data types and knowledge representations within the neuroscience domain, the resulting models exhibit strong generalization and broad applicability.
@@ 3 Methodology
Figure 1: The pipeline of our proposed BrainGFM. (a) A large-scale brain fMRI graph dataset is constructed for pre-training. (b) BrainGFM is pre-trained using graph contrastive and masked autoencoder strategies, with atlas/parcellation tokens [A/P] to encode atlas-specific information. (c) We introduce graph prompts and use meta-learning to optimize them for few-shot adaptation, keeping the graph FM backbone frozen. (d) Finally, we freeze both the model and graph prompts, and use language prompts to enable zero-shot transfer to new tasks. Note that ”Schf.” means Schaefer atlas.
@@ 3.2 Graph Pre-Training for Building BrainGFM
We adopt a Graph Transformer (Yun et al., 2019) as the backbone of our BrainGFM. As illustrated in Figure 1(b), we first transform the input fMRI brain graphs and project them to obtain brain graph embeddings, where each token corresponds to a brain ROI. We apply Positional Encoding Tokens, denoted as [PE], to each brain ROI, enabling the model to perceive and learn the topological and spatial characteristics of each ROI. Unlike conventional positional encodings used in standard Transformer models (Vaswani et al., 2017), graph-based positional encodings are inherently different, as they require encoding the relative positions between nodes in the graph structure. Compared to the commonly used Laplacian positional encoding (Dwivedi et al., 2023) and node degree positional encoding (You et al., 2019) in graph-based models, we adopt a more efficient alternative Random Walk Structural Encoding (RWSE) (Dwivedi et al., 2021) as our positional encoding strategy. More details about PEs can be found in Appendix I. Furthermore, inspired by language models in NLP, we insert Atlas/Parcellation Tokens, denoted as [A/P] and Task/Disorder Tokens, denoted as [T/D], to the brain graph embeddings during pre-training, enabling the model to better distinguish and learn from different atlases and parcellations. Note that the construction of [A/P] and [T/D] tokens is described in detail in Section 3.4. Incorporating this token enables the model to capture parcellation-specific patterns, which is crucial as prior studies (Hermosillo et al., 2024; Liu et al., 2023; Wu et al., 2025) show that different brain disorders are better represented by specific parcellations. For instance, MDD benefits from Schaefer200 or Power264 and ASD is better captured by Shen268 or Schaefer200. Embedding such parcellation-aware information helps improve model generalization across disorders and atlas.
@@ 3.4 Zero-Shot Graph/Language Prompt-Tuning
Building on the few-shot capability, we further introduce language prompts to enable more generalized zero-shot learning by jointly guiding both the graph prompt and the pre-trained foundation model. In zero-shot scenarios, the parameters of the graph prompt are also frozen, meaning the model cannot rely on prompt adaptation through learning. Instead, the language prompt provides semantic guidance, allowing the model to generalize and adapt to unseen downstream data, tasks, and disorder types without any gradient-based updates. As shown in Figure 1(d), in order to enable the model to recognize and distinguish between different tasks and disorders in zero-shot settings, we introduce Task/Disorder Tokens, denoted as [T/D]. To construct the [T/D] tokens, we first generate a textual description for each disorder, including its full name, abbreviation, and a concise clinical summary. For example, for Major Depressive Disorder, the corresponding text description is: Major Depressive Disorder (MDD) is a common mental illness characterized by persistent and profound low mood, loss of interest, and cognitive impairment, significantly affecting daily life and social functioning.” (Otte et al., 2016) We then encode these textual descriptions using a BERT model (Devlin et al., 2019) pre-trained on large-scale medical corpora, such as BioClinicalBERT (Huang et al., 2019; Alsentzer et al., 2019), to obtain semantic-rich text embeddings. These embeddings are subsequently projected and embedded as [T/D] tokens, which are incorporated into the model during downstream adaptation. Similarly, the construction of [A/P] tokens is also based on language text. For each atlas and parcellation, we provide a textual description of its name, such as “Schaefer100”, “Schaefer200”, or “AAL116”. These text descriptions are then encoded using the BioClinicalBERT pre-trained model to extract language embeddings, which are subsequently transformed into [A/P] tokens. As shown in Figure 1(d), the [T/D] and [A/P] tokens are ultimately concatenated with the ROI tokens from the graph embeddings as language prompt tokens. This combined input is then fed into the foundation model to guide feature extraction specific to the given dataset, task, and disorder. By introducing disorder-specific semantic priors through the [T/D] and [A/P] tokens, the model is better equipped to capture characteristics from different tasks, disorders, atlas and parcellations, thereby improving its downstream adaptation ability in zero-shot settings without any training.
@@ 4.1 Comparison with Other Methods
BrainLM 67.6±1.9 69.1±2.2 64.0±2.7 71.7±2.0 68.1±2.6 71.6±1.8 66.9±2.5 69.6±2.4 78.3±1.7 79.7±2.5 73.1±2.6 81.4±2.7 77.8±2.0 80.7±2.1 81.3±2.8 77.9±2.4 82.6±2.5 85.0±2.6 78.6±2.0 86.5±2.7
BrainMass 67.0±2.3 67.5±2.0 64.7±2.8 71.1±1.8 68.9±2.1 70.1±1.9 69.5±2.4 66.3±2.6 77.8±2.3 80.6±2.7 72.6±2.9 81.4±2.5 76.9±2.4 81.5±2.8 80.6±2.5 78.1±2.6 81.0±2.7 84.0±2.2 80.3±2.3 81.9±2.9
Brain-JEPA 69.8±1.9 71.6±2.0 66.2±1.6 72.9±2.4 70.1±1.7 73.8±1.8 70.1±1.9 69.3±1.7 79.1±2.2 81.6±1.8 76.8±2.2 83.6±2.1 79.3±1.7 82.8±1.3 82.3±1.9 78.5±1.5 85.4±1.8 86.7±2.0 85.3±2.2 83.7±1.9
BrainGFM 70.3±1.6 70.5±1.5 67.0±1.7 71.4±2.1 71.2±1.9 73.5±1.4 70.4±1.5 69.8±1.7 80.3±2.6 81.9±2.2 76.2±1.5 84.4±1.7 79.9±1.6 82.0±1.7 83.2±1.6 79.2±1.9 85.2±1.6 86.3±2.1 87.7±1.9 82.6±1.7
@@ 4.2 Ablation Study for Full/Few/Zero-Shot on Graph/Language Prompt and Meta Learning
4.3 Experiments on Different Atlases and Parcellations
@@ 4.4 Comparison Among Time-Series, ROI and Graph-Based Foundation Models
Figure 4: Comparison of performance and efficiency across different FMs.
@@ 4.6 Ablation Study on Different Foundation Model Pre-Training Methods
As shown in Figure 5, we compare the effectiveness of different graph pre-training strategies, including graph contrastive learning, graph-masked autoencoders and their sequential combination. The results demonstrate that GCL slightly outperforms GMAE, and that combining GCL and GMAE yields further performance gains compared to using either method alone.
@@
GCL pre-training primarily focuses on capturing global representations of brain graphs by encouraging the model to aggregate holistic graph-level features and distinguish between different graph attributes and categories. In comparison, GMAE pre-training emphasizes the learning of local representations, where the model reconstructs masked brain ROIs based on information from their local neighborhoods, thereby promoting specialization in ROI-level feature extraction. By sequentially combining GCL and GMAE during pre-training, BrainGFM is able to simultaneously acquire both global and local discriminative capabilities. Notably, the integration of global and local information has been widely recognized as critical for understanding brain organization and pathology in neuroscience and neuroimaging studies. Consequently, our pre-trained model benefits from this multi-scale representation learning, leading to enhanced transferability and improved performance across various downstream tasks.
5 Broader Impact
6 Limitation and Future Works
While our work successfully constructed a large-scale fMRI dataset for pre-training, certain limitations remain. Due to the significant manual effort involved, we were unable to include all datasets from the OpenNeuro platform (Markiewicz et al., 2021), particularly the large number of task-based (non-resting-state) fMRI datasets. In addition, because of financial constraints, we were not able to incorporate fMRI data from the UK Biobank (Bycroft et al., 2018), including both resting-state and task-based scans, as access to this dataset requires paid licensing and ongoing maintenance costs.
@@ Appendix E Ablation Study for Different Brain Graph Construction Methods
Table 3: Comparison of different brain graph construction methods on HBN dataset for MDD diagnosis. Our approach uses top-$k$ sparsification of the Pearson correlation matrix to construct brain graphs. We also compare it with KNN-based graph construction. The numbers in the table represent classification accuracy.
Corpus Pearson Top-K KNN Partial Correlation Mutual Information Dynamic Connectivity HBN MDD 85.5 86.7 84.8 83.9 85.7
@@ Appendix G Ablation Study on Different Tuning Methods
As shown in Table 5, after completing model pre-training, we explored three downstream adaptation strategies: full fine-tuning (Sun et al., 2022), parameter-efficient fine-tuning (PEFT) (Ding et al., 2023), and graph prompt-tuning (Sun et al., 2023). Full fine-tuning updates all model parameters during downstream training, offering strong performance but at a high computational cost. In contrast, PEFT methods reduce training overhead by modifying only a small subset of parameters or introducing lightweight modules. We specifically evaluate two popular PEFT variants: prefix-tuning and LoRA. Note that the details of all tuning methods are summarized in Table 6.
Graph prompt-tuning further improves efficiency by freezing the entire pre-trained model and updating only a small set of learnable prompt vectors injected into the input. This strategy allows the model to adapt without altering its core parameters, making it highly suitable for resource-constrained settings. In full-shot scenarios, fine-tuning delivers the best results, but PEFT methods achieve competitive performance with significantly lower computational demands. Graph prompt-tuning, while slightly less accurate, offers the best efficiency–adaptability trade-off by minimizing trainable parameters. Given the structural nature of brain graphs, where node and edge features capture complex spatial and relational dependencies, we examine two prompt insertion mechanisms: addition (“+”) and multiplication ($\times$”). Results show that multiplicative insertion consistently outperforms the additive version, likely because scaling features better preserves relational patterns. Moreover, to further leverage topological information, we extend the prompt-tuning framework by incorporating edge prompts in addition to node prompts. This design grants the model greater flexibility to adjust local connectivity, leading to improved transfer performance across diverse downstream tasks.